-solve using long division
(x^3-64)divided by (x-4)
please show me how to do this problem and then a check

Answer by sachi(548)

x^3-64=x^3-4^3=(x-4)(x^2+4x+4^2)
so (x^3-64)/(x-4)=x^2+4x+4^2
by long division
x-4)x^3-64(x^2+4x+16 (ans)
x^3-4x^2
- +
--------------
4x^2-64
4x^2-16x
- +
--------------
16x-64
16x-64
- +
--------------
0